# On the quantitative lower bounds for solutions to the Boltzmann equation in nonconvex domains

### v3

@@ 4. Maxwellian bound for the cut-off case for non-fully specular reflection condition
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the collision kernel $B$ satisfies Assumption B with $\alpha\in[0,1]$. We consider a continuous mild solution $f(t,x,v)$ of (1.1)–(1.3) . Given $\tau^{\prime\prime}\in(0,\Delta^{0})$ with $\Delta^{0}>0$ given in Proposition 3.1, there exist $b(\tau^{\prime\prime})$, $\delta_{T,I}(\tau^{\prime\prime})>0$, which depends on $\Omega,M,E_{f}$, and the modulus of continuity of $f$, such that for $t\in[\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T,I}(\tau^{\prime\prime})]$,
@@
Then, using Proposition 3.2 (note that as before by (LABEL:2025_09/09_06:47), the assumption (3.9) holds with $\tau=\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})$, $\Delta_{1}=2\mathfrak{d}(\tau^{\prime\prime}),\ \Delta_{2}=\mathfrak{d}(\tau^{\prime\prime})$, $x^{\prime}=y^{0}_{i}$, and $v^{\prime}=v_{i}(\tau^{\prime\prime})$, we also have $B(y^{0}_{i},\mathfrak{d}(\tau^{\prime\prime}))\subset B(y^{0}_{i},\frac{d}{8})\subset\Omega$ by Lemma 2.2), we have for $\tau^{\prime\prime}\leq t\leq T_{\partial}+\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}+1}\left(\frac{2d_{\Omega}}{T_{\partial}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}+1\right)}$ (notice Remark 7 and the fact that $T_{\partial}+\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}+1}\left(\frac{2d_{\Omega}}{T_{\partial}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}+1\right)}\leq\tau^{\prime\prime}+\mathfrak{d}(\tau^{\prime\prime})$)
@@
(4.14) $\begin{split}&f(t,x_{\partial},v)\\ \geq&f\left(\frac{t+(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2},x_{\partial}-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v,v\right)e^{-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}C_{L}\left<R^{\partial}_{\min}(\tau^{\prime\prime})+1\right>^{\gamma^{+}}}\\ \geq&\alpha_{\mathfrak{m}_{\partial}}\left(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime})\right)\\ &\mathbf{1}_{B\left(y^{0}_{i},\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}}}\right)\times B(v_{i}(\tau^{\prime\prime}),r_{\mathfrak{m}_{\partial}}(\mathfrak{d}(\tau^{\prime\prime})))}\left(x_{\partial}-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v,v\right)e^{-C_{L}\left<R^{\partial}_{\min}(\tau^{\prime\prime})+1\right>^{\gamma^{+}}}.\end{split}$
@@
(4.16) $\begin{split}&\alpha_{\mathfrak{m}_{\partial}}(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\frac{2\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}))\\ &\mathbf{1}_{B\left(y^{0}_{i},\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}}}\right)\times B(v_{i}(\tau^{\prime\prime}),r_{\mathfrak{m}_{\partial}}(\mathfrak{d}(\tau^{\prime\prime})))}(x_{\partial}-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v,v)e^{-C_{L}\left<R^{\partial}_{\min}(\tau^{\prime\prime})+1\right>^{\gamma^{+}}}\\ \geq&\alpha_{\mathfrak{m}_{\partial}}(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\frac{2\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}))e^{-C_{L}\left<R^{\partial}_{\min}(\tau^{\prime\prime})+1\right>^{\gamma^{+}}}.\end{split}$
@@
$\geq$ $\alpha_{\mathfrak{m}_{\partial}}(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\frac{2\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}))e^{-C_{L}\left<R^{\partial}_{\min}(\tau^{\prime\prime})+1\right>^{\gamma^{+}}}$
@@
$\geq$ $\alpha_{\mathfrak{m}_{\partial}}(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\frac{2\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}))e^{-C_{L}\left<R^{\partial}_{\min}(\tau^{\prime\prime})+1\right>^{\gamma^{+}}}\mathbf{1}_{B(v_{x_{\partial}}(\tau^{\prime\prime}),\frac{1}{2^{\mathfrak{m}_{\partial}+2}})}(v).$
@@
$A^{\prime}:=\alpha_{\mathfrak{m}_{\partial}}(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\frac{2\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}))e^{-C_{L}\left<R^{\partial}_{\min}(\tau^{\prime\prime})+1\right>^{\gamma^{+}}},$
@@
The rest of the proof of existence of lower bound is exactly the same as in Briant’s paper [1](From Prop 3.9, p21 p25). Before we introduce the next proposition, we define some constants here:
@@
(4.22) $b_{n,\tau^{\prime\prime}}(\tau^{\prime}):=\begin{cases}\min\{b(\tau^{\prime\prime}),a_{0}(\tau^{\prime\prime})\}&,\ n=0\\ \min\begin{Bmatrix}(1-\alpha)b(\tau^{\prime\prime})e^{-C_{L}(\tau^{\prime}-\tau^{\prime\prime})\langle\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{1}\rangle^{\gamma^{+}}}\\ \times\frac{1}{2\pi T_{B}^{2}}e^{-\frac{(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{1})^{2}}{2T_{B}}},\\ (\min(a_{0}(\tau^{\prime\prime}),b(\tau^{\prime})))^{2}C_{Q}r_{0}^{{}^{\prime\prime}3+\gamma}\xi^{\frac{1}{2}}_{1}\\ \times\frac{\tau^{\prime}-\tau^{\prime\prime}}{2^{2}(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{1})}e^{-C_{L}\frac{(\tau^{\prime}-\tau^{\prime\prime})\langle\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{1}\rangle^{\gamma^{+}}}{2(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{1})}}\end{Bmatrix}&,\ n=1\\ (1-\alpha)b(\tau^{\prime\prime})e^{-C_{L}(\tau^{\prime}-\tau)\langle\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n}\rangle^{\gamma^{+}}}\frac{1}{2\pi T_{B}^{2}}e^{-\frac{(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n})^{2}}{2T_{B}}}&,\ n\geq 2,\end{cases}$
where $b(\tau_{0})$ is introduced in Prop. 4.1, and
@@
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the collision kernel $B$ satisfies Assumption B with $\nu<0$. Let $f$ be the continuous mild solution of (1.1)–(1.3) with $0\leq\alpha<1$. We fix $\tau^{\prime\prime}\in(0,\Delta^{0})$, where $\Delta^{0}>0$ is given in Proposition 3.1. Then, there exist $\delta_{T,D}(\tau^{\prime\prime})>0$, which depends on $\Omega,M,E_{f}$, and the modulus of continuity of $f$, such that for every $\tau^{\prime}\in(\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T,D}(\tau^{\prime\prime})]$ and every $N\in\mathbb{N}$, for each $0\leq n\leq N$ and each $1\leq i\leq\mathcal{N}(N,\tau^{\prime\prime})$, we have
@@
First, we observe that by definition of $t_{\partial}$ in Definition 1.1, we have $t_{\partial}\leq t-s\leq t-\tau^{\prime\prime}\leq\tau^{\prime}-\tau^{\prime\prime}$. Next, we notice by (4.25) that $\tau^{\prime}\in(\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T,I}(\tau^{\prime\prime})]$. Then, since $\alpha<1$, the diffuse part yields a strictly positive contribution, thus we use (1.13), Proposition 4.1, and Lemma 3.1 to derive that for any $(t,x,v)$ satisfying (4.26), we have
(4.36) $\begin{split}&f(t,x,v)\\ \geq&(1-\alpha)\left(\int_{w\cdot n(X_{t-t_{\partial},t}(x,v))>0}f(t,X_{t-t_{\partial},t}(x,v),w)(w\cdot n(X_{t-t_{\partial},t}(x,v)))dw\right)\\ &\frac{1}{2\pi T_{B}^{2}}e^{-\frac{|v|^{2}}{2T_{B}}}\exp\left[-\int_{t-t_{\partial}}^{t}L[f(s,X_{s,t}(x,v),\cdot)](v)ds\right]\\ \geq&(1-\alpha)b(\tau^{\prime\prime})\frac{1}{2\pi T_{B}^{2}}e^{-\frac{|v|^{2}}{2T_{B}}}e^{-t_{\partial}C_{L}\left\langle v\right\rangle^{\gamma^{+}}}\\ \geq&(1-\alpha)b(\tau^{\prime\prime})\frac{1}{2\pi T_{B}^{2}}e^{-\frac{(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{k+1})^{2}}{2T_{B}}}e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}{\left\langle\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{k+1}\right\rangle}^{\gamma^{+}}}\\ \geq&b_{k+1,\tau^{\prime\prime}}(\tau^{\prime})\geq\min\left(a_{k+1,\tau^{\prime\prime}}(\tau^{\prime}),b_{k+1,\tau^{\prime\prime}}(\tau^{\prime})\right),\end{split}$
@@
Finally, we end the proof of Theorem 1.1 for the case $\alpha\in[0,1)$ by picking $0<\xi<\frac{1}{4}$ and setting $\xi_{n}:=\xi^{n}$, by which we can deduce that (we recall $r_{n}^{\prime\prime}$ from (4.21))
@@
Notice first that it suffices to prove that given $\tau_{2}\in(\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T,D}(\tau^{\prime\prime})]$ there exists $\mathcal{A}>0$ such that
@@
$\geq$ $(1-\alpha)b(\tau^{\prime\prime})\frac{1}{2\pi T_{B}^{2}}e^{-8C_{L}\tau_{2}\mathfrak{R}_{\min}(\tau^{\prime\prime})^{\gamma^{+}}}e^{-\frac{\mathfrak{R}_{\min}(\tau^{\prime\prime})^{2}}{T_{B}}}e^{-\left[(2C_{L}\tau_{2}\mathfrak{R}_{\min}(\tau^{\prime\prime})^{\gamma^{+}}+\frac{1}{T_{B}})(\delta_{V}(\tau^{\prime\prime}))^{2}2^{n}\right]}$
$\geq$ $C_{1}e^{-\left[(2C_{L}\tau_{2}\mathfrak{R}_{\min}(\tau^{\prime\prime})^{\gamma^{+}}+\frac{1}{T_{B}})(\delta_{V}(\tau^{\prime\prime}))^{2}2^{n}\right]},$
@@
$\mathcal{A}_{1}(\tau_{2}):=\begin{cases}\min\left\{b_{0,\tau^{\prime\prime}}(\tau_{2}),\sqrt{b_{1,\tau^{\prime\prime}}(\tau_{2})},e^{-\left[(2C_{L}\tau_{2}\mathfrak{R}_{\min}(\tau^{\prime\prime})^{\gamma^{+}}+\frac{1}{T_{B}})(\delta_{V}(\tau^{\prime\prime}))^{2}\right]}\right\},\ C_{1}\geq 1\\ \min\left\{b_{0,\tau^{\prime\prime}}(\tau_{2}),\sqrt{b_{1,\tau^{\prime\prime}}(\tau_{2})},C_{1}e^{-\left[(2C_{L}\tau_{2}\mathfrak{R}_{\min}(\tau^{\prime\prime})^{\gamma^{+}}+\frac{1}{T_{B}})(\delta_{V}(\tau^{\prime\prime}))^{2}\right]}\right\},\ C_{1}<1,\end{cases}$
@@
$C_{2}:=C_{Q}(\tau_{2}-\tau^{\prime\prime})e^{-C_{L}\left[\frac{\tau_{2}(1+\mathfrak{R}_{\min}(\tau^{\prime\prime}))^{\gamma^{+}}}{\mathfrak{R}_{\min}(\tau^{\prime\prime})}+\frac{\tau_{2}(\delta_{V}(\tau^{\prime\prime}))^{\gamma^{+}}}{\mathfrak{R}_{\min}(\tau^{\prime\prime})}\right]},$

### v4

@@ 4. Maxwellian bound for the cut-off case for non-fully specular reflection condition
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the collision kernel $B$ satisfies Assumption B with $\alpha\in[0,1]$. We consider a continuous mild solution $f(t,x,v)$ of (1.1)–(1.3) . Given $\tau^{\prime\prime}\in(0,\Delta^{0})$ with $\Delta^{0}>0$ given in Proposition 3.1, there exist $b(\tau^{\prime\prime})$, $\delta_{T,I}(\tau^{\prime\prime})>0$, which depends on $\Omega,M,E_{f}$, and the modulus of continuity of $f_{0}$, such that for $t\in[\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T,I}(\tau^{\prime\prime})]$,
@@
(4.2) $f(t,x,v)\geq\ \mathfrak{B}(\tau^{\prime\prime})\mathbf{1}_{B(y^{0}_{i},\mathfrak{d}(\tau^{\prime\prime}))\times B(v_{i}(\tau^{\prime\prime}),\mathfrak{d}(\tau^{\prime\prime}))}(x,v),\ \forall t\in[\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}+\mathfrak{d}(\tau^{\prime\prime})],$
@@
Then, using Proposition 3.2 (note that as before by (4.2), the assumption (3.9) holds with $\tau=\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})$, $\Delta_{1}=2\mathfrak{d}(\tau^{\prime\prime}),\ \Delta_{2}=\mathfrak{d}(\tau^{\prime\prime})$, $x^{\prime}=y^{0}_{i}$, and $v^{\prime}=v_{i}(\tau^{\prime\prime})$, we also have $B(y^{0}_{i},\mathfrak{d}(\tau^{\prime\prime}))\subset B(y^{0}_{i},\frac{d}{8})\subset\Omega$ by Lemma 2.2), we have for $\tau^{\prime\prime}\leq t\leq T_{\partial}+\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}+1}\left(\frac{2d_{\Omega}}{T_{\partial}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}+1\right)}$ (see Remark 3.2 and the fact that $T_{\partial}+\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}+1}\left(\frac{2d_{\Omega}}{T_{\partial}-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}+1\right)}\leq\tau^{\prime\prime}+\mathfrak{d}(\tau^{\prime\prime})$)
@@
(4.14) $\begin{split}&f(t,x_{\partial},v)\\ \geq&f\left(\frac{t+(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2},x_{\partial}-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v,v\right)e^{-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}C_{L}\left\langle R^{\partial}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}\\ \geq&\alpha_{\mathfrak{m}_{\partial}}\left(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\tau^{\prime\prime}-\frac{\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime})\right)\\ &\mathbf{1}_{B\left(y^{0}_{i},\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}}}\right)\times B(v_{i}(\tau^{\prime\prime}),r_{\mathfrak{m}_{\partial}}(\mathfrak{d}(\tau^{\prime\prime})))}\left(x_{\partial}-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v,v\right)e^{-C_{L}\left\langle R^{\partial}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}.\end{split}$
@@
(4.16) $\begin{split}&\alpha_{\mathfrak{m}_{\partial}}(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\frac{2\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}))\\ &\mathbf{1}_{B\left(y^{0}_{i},\frac{\mathfrak{d}(\tau^{\prime\prime})}{2^{\mathfrak{m}_{\partial}}}\right)\times B(v_{i}(\tau^{\prime\prime}),r_{\mathfrak{m}_{\partial}}(\mathfrak{d}(\tau^{\prime\prime})))}(x_{\partial}-\frac{t-(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}))}{2}v,v)e^{-C_{L}\left\langle R^{\partial}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}\\ \geq&\alpha_{\mathfrak{m}_{\partial}}(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\frac{2\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}))e^{-C_{L}\left\langle R^{\partial}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}.\end{split}$
@@
$\geq$ $\alpha_{\mathfrak{m}_{\partial}}(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\frac{2\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}))e^{-C_{L}\left\langle R^{\partial}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}$
@@
$\geq$ $\alpha_{\mathfrak{m}_{\partial}}(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\frac{2\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}))e^{-C_{L}\left\langle R^{\partial}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}}\mathbf{1}_{B(v_{x_{\partial}}(\tau^{\prime\prime}),\frac{1}{2^{\mathfrak{m}_{\partial}+2}})}(v).$
@@
$A^{\prime}:=\alpha_{\mathfrak{m}_{\partial}}(\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime}),\frac{2\tau^{\prime\prime}-\mathfrak{d}(\tau^{\prime\prime})}{2},\mathfrak{d}(\tau^{\prime\prime}),\mathfrak{B}(\tau^{\prime\prime}),\mathfrak{R}(\tau^{\prime\prime}))e^{-C_{L}\left\langle R^{\partial}_{\min}(\tau^{\prime\prime})+1\right\rangle^{\gamma^{+}}},$
@@
The rest of the proof of existence of lower bound is exactly the same as in Briant’s paper [1](from Prop. 3.9, pp. 21–25). Before we introduce the next proposition, we define some constants here:
@@
(4.22) $b_{n,\tau^{\prime\prime}}(\tau^{\prime}):=\begin{cases}\min\{b(\tau^{\prime\prime}),a_{0}(\tau^{\prime\prime})\}&,\ n=0\\ \min\begin{Bmatrix}(1-\alpha)b(\tau^{\prime\prime})e^{-C_{L}(\tau^{\prime}-\tau^{\prime\prime})\langle\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{1}\rangle^{\gamma^{+}}}\\ \times\frac{1}{2\pi T_{B}^{2}}e^{-\frac{(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{1})^{2}}{2T_{B}}},\\ (\min(a_{0}(\tau^{\prime\prime}),b(\tau^{\prime})))^{2}C_{Q}r_{0}^{{}^{\prime\prime}3+\gamma}\xi^{\frac{1}{2}}_{1}\\ \times\frac{\tau^{\prime}-\tau^{\prime\prime}}{2^{2}(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{1})}e^{-C_{L}\frac{(\tau^{\prime}-\tau^{\prime\prime})\langle\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{1}\rangle^{\gamma^{+}}}{2(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{1})}}\end{Bmatrix}&,\ n=1\\ (1-\alpha)b(\tau^{\prime\prime})e^{-C_{L}(\tau^{\prime}-\tau^{\prime\prime})\langle\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n}\rangle^{\gamma^{+}}}\frac{1}{2\pi T_{B}^{2}}e^{-\frac{(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{n})^{2}}{2T_{B}}}&,\ n\geq 2,\end{cases}$
where $b(\tau^{\prime\prime})$ is introduced in Prop. 4.1, and
@@
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the collision kernel $B$ satisfies Assumption B with $\nu<0$. Let $f$ be the continuous mild solution of (1.1)–(1.3) with $0\leq\alpha<1$. We fix $\tau^{\prime\prime}\in(0,\Delta^{0})$, where $\Delta^{0}>0$ is given in Proposition 3.1. Then, there exist $\delta_{T,D}(\tau^{\prime\prime})>0$, which depends on $\Omega,M,E_{f}$, and the modulus of continuity of $f_{0}$, such that for every $\tau^{\prime}\in(\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T,D}(\tau^{\prime\prime})]$ and every $N\in\mathbb{N}$, for each $0\leq n\leq N$ and each $1\leq i\leq\mathcal{N}(N,\tau^{\prime\prime})$, we have
@@
First, we observe that by definition of $t_{\partial}(x,v)$ in Definition 1.1, we have $t_{\partial}(x,v)\leq t-s\leq t-\tau^{\prime\prime}\leq\tau^{\prime}-\tau^{\prime\prime}$. Next, we notice by (4.25) that $\tau^{\prime}\in(\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T,I}(\tau^{\prime\prime})]$. Then, since $\alpha<1$, the diffuse part yields a strictly positive contribution, thus we use (1.13), Proposition 4.1, and Lemma 3.1 to derive that for any $(t,x,v)$ satisfying (4.26), we have
(4.36) $\begin{split}&f(t,x,v)\\ \geq&(1-\alpha)\left(\int_{w\cdot n(X_{t-t_{\partial}(x,v),t}(x,v))>0}f(t,X_{t-t_{\partial}(x,v),t}(x,v),w)(w\cdot n(X_{t-t_{\partial}(x,v),t}(x,v)))dw\right)\\ &\frac{1}{2\pi T_{B}^{2}}e^{-\frac{|v|^{2}}{2T_{B}}}\exp\left[-\int_{t-t_{\partial}(x,v)}^{t}L[f(s,X_{s,t}(x,v),\cdot)](v)\,ds\right]\\ \geq&(1-\alpha)b(\tau^{\prime\prime})\frac{1}{2\pi T_{B}^{2}}e^{-\frac{|v|^{2}}{2T_{B}}}e^{-t_{\partial}(x,v)C_{L}\left\langle v\right\rangle^{\gamma^{+}}}\\ \geq&(1-\alpha)b(\tau^{\prime\prime})\frac{1}{2\pi T_{B}^{2}}e^{-\frac{(\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{k+1})^{2}}{2T_{B}}}e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}{\left\langle\mathfrak{R}_{\min}(\tau^{\prime\prime})+r^{\prime\prime}_{k+1}\right\rangle}^{\gamma^{+}}}\\ \geq&b_{k+1,\tau^{\prime\prime}}(\tau^{\prime})\geq\min\left(a_{k+1,\tau^{\prime\prime}}(\tau^{\prime}),b_{k+1,\tau^{\prime\prime}}(\tau^{\prime})\right),\end{split}$
@@
Finally, we end the proof of Theorem 1.1 for case $\alpha\in[0,1)$. We pick $0<\xi<\frac{1}{4}$ and setting $\xi_{n}:=\xi^{n}$, by which we can deduce that (we recall $r_{n}^{\prime\prime}$ from (4.21))
@@
We first show that given $\tau_{2}\in\left[\tau^{\prime\prime}+\min\left\{\frac{\delta_{T,D}(\tau^{\prime\prime})-\tau^{\prime\prime}}{2},\tau^{\prime\prime}\right\},\tau^{\prime\prime}+\delta_{T,D}(\tau^{\prime\prime})\right]$ there exists $\mathcal{A}>0$ such that
@@
$\geq$ $(1-\alpha)b(\tau^{\prime\prime})\frac{1}{2\pi T_{B}^{2}}e^{-8C_{L}\tau_{2}\mathfrak{R}_{\min}(\tau^{\prime\prime})^{\gamma^{+}}}e^{-\frac{\mathfrak{R}_{\min}(\tau^{\prime\prime})^{2}}{T_{B}}}e^{-\left((2C_{L}\tau_{2}\mathfrak{R}_{\min}(\tau^{\prime\prime})^{\gamma^{+}}+\frac{1}{T_{B}})(\delta_{V}(\tau^{\prime\prime}))^{2}2^{n}\right)}$
$\geq$ $C_{1}e^{-\left((2C_{L}\tau_{2}\mathfrak{R}_{\min}(\tau^{\prime\prime})^{\gamma^{+}}+\frac{1}{T_{B}})(\delta_{V}(\tau^{\prime\prime}))^{2}2^{n}\right)},$
@@
$\mathcal{A}_{1}(\tau_{2}):=\begin{cases}\min\left\{b_{0,\tau^{\prime\prime}}(\tau_{2}),\sqrt{b_{1,\tau^{\prime\prime}}(\tau_{2})},e^{-\left((2C_{L}\tau_{2}\mathfrak{R}_{\min}(\tau^{\prime\prime})^{\gamma^{+}}+\frac{1}{T_{B}})(\delta_{V}(\tau^{\prime\prime}))^{2}\right)}\right\},\ C_{1}\geq 1\\ \min\left\{b_{0,\tau^{\prime\prime}}(\tau_{2}),\sqrt{b_{1,\tau^{\prime\prime}}(\tau_{2})},C_{1}e^{-\left((2C_{L}\tau_{2}\mathfrak{R}_{\min}(\tau^{\prime\prime})^{\gamma^{+}}+\frac{1}{T_{B}})(\delta_{V}(\tau^{\prime\prime}))^{2}\right)}\right\},\ C_{1}<1,\end{cases}$
@@
$C_{2}:=C_{Q}\min\left\{\frac{\delta_{T,D}(\tau^{\prime\prime})-\tau^{\prime\prime}}{2},\tau^{\prime\prime}\right\}e^{-C_{L}\left(\frac{\tau_{2}(1+\mathfrak{R}_{\min}(\tau^{\prime\prime}))^{\gamma^{+}}}{\mathfrak{R}_{\min}(\tau^{\prime\prime})}+\frac{\tau_{2}(\delta_{V}(\tau^{\prime\prime}))^{\gamma^{+}}}{\mathfrak{R}_{\min}(\tau^{\prime\prime})}\right)},$